### v4

# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

Abstract
Vision-language models (VLMs) have achieved impressive progress in natural image reasoning, yet their potential in medical imaging remains underexplored. Medical vision-language tasks demand precise understanding and clinically coherent answers, which are difficult to achieve due to the complexity of medical data and the scarcity of high-quality expert annotations. These challenges limit the effectiveness of conventional supervised fine-tuning (SFT) and Chain-of-Thought (CoT) strategies that work well in general domains. To address these challenges, we propose Med-R1, a reinforcement learning (RL)-enhanced vision-language model designed to improve generalization and reliability in medical reasoning. Built on the DeepSeek strategy, Med-R1 adopts Group Relative Policy Optimization (GRPO) to encourage reward-guided learning beyond static annotations. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities. Med-R1 achieves a 29.94% improvement in average accuracy over its base model Qwen2-VL-2B, and even outperforms Qwen2-VL-72B—a model with 36$\times$ more parameters. To assess cross-task generalization, we further evaluate Med-R1 on five question types. Med-R1 outperforms Qwen2-VL-2B by 32.06% in question-type generalization, also surpassing Qwen2-VL-72B. We further explore the thinking process in Med-R1, a crucial component for the success of Deepseek-R1. Our results show that omitting intermediate rationales (No-Thinking-Med-R1) not only improves in-domain and cross-domain generalization with less training, but also challenges the assumption that more reasoning always helps. These findings suggest that in medical VQA, it is not reasoning itself, but its quality and domain alignment, that determine effectiveness. Together, these results highlight that RL improves medical reasoning and generalization, enabling efficient and reliable VLMs for real-world deployment.
{IEEEkeywords}
Reinforcement Learning, Vision Language Models, Multimodal LLM, Post-Training, Medical Reasoning.
1 Introduction
[…]
Vision-language models (VLMs) have achieved strong performance on natural image understanding tasks such as visual question answering (VQA) and multimodal dialogue [1, 2, 3], enabled by large-scale pretraining and supervised fine-tuning (SFT). However, applying VLMs to medical imaging remains challenging due to the need for clinically sound interpretations and decision-making processes. Medical tasks often involve multi-step analysis—e.g., diagnosing a lung nodule may require integrating lesion localization, morphology, and context. In addition, the diversity of imaging modalities (e.g., CT, MRI) and task types (e.g., diagnosis, grading) imposes demands on generalizability. This raises a key question: how can we enable VLMs to perform well across medical domains while ensuring reliable and context-aware behavior?
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [4, 5]. While SFT has been widely adopted to adapt foundation models to medical imaging [6, 7, 8, 9, 10], it suffers from two fundamental issues that hinder medical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [6, 7] lack access to high-quality CoT data, leading to shallow reasoning with limited clinical rigor. These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks. This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are indispensable requirements.
To address these challenges, we propose Med-R1, a reinforcement learning (RL)-based framework for enhancing the generalizability and interpretability of medical VLMs. Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11]. We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO [13] that stabilizes training via rule-based rewards and group-relative comparisons. These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
As shown in Figure 1, we evaluate Med-R1 across eight diverse medical imaging modalities, including CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy, and X-ray. These modalities cover a wide range of clinical imaging—from macroscopic anatomy to cellular-level and functional assessments. Med-R1 (2B parameters) achieves 69.91% average accuracy, a 29.94% gain over its base model Qwen2-VL-2B, and even outperforms the 72B-parameter Qwen2-VL-72B (Table 3), highlighting the benefit of RL-driven adaptation. We further assess generalizability across five question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 6). These demonstrate that RL enhances both parameter efficiency and task generalizability in medical VLMs.
Finally, inspired by recent findings on reasoning behavior in rule-based RL for classification tasks [14], we further investigate the role of intermediate rationales in Med-R1. A widely held assumption in VLM development is that generating step-by-step reasoning (“Think”) enhances generalization and performance. However, our empirical analysis reveals that this may not hold in domain-specific settings like medical VQA. We compare two RL post-training strategies: one that guides the model to produce explicit rationales before answering, and another that directly outputs the final answer without any intermediate reasoning steps. Surprisingly, we observe that the latter (No-Thinking) strategy often yields better generalization across both imaging modalities and question types. We hypothesize that in the absence of high-quality CoT supervision, RL-trained models fall back on their pre-trained reasoning patterns, which are typically tuned for general-domain tasks. Under large domain shifts—such as differences in image characteristics, task goals, and clinical language—this can result in hallucinated or misaligned rationales that impair performance. These findings suggest that it is not the presence of reasoning per se, but rather its quality and domain alignment, that determine its utility in medical vision-language reasoning.
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
  1. 1.
    Multi-modality multi-task reasoning medical VLM. We propose Med-R1, the first reasoning medical VLM capable of supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.) across five unique tasks. We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT. The proposed Med-R1 is capable of generating step-by-step, accurate, and clinically plausible explanations.
  2. 2.
    Robust Generalization with Efficiency. We show that alongside the solid modality and task-specific performance, Med-R1 exhibits strong generalization. Med-R1 outperforms the base model by 29.94% and SFT baselines by 15.84% in average generalization accuracy across modalities. In cross-task settings, it outperforms the base model and SFT baseline by 32.06% and 11.25% respectively. Moreover, Med-R1 surpasses other larger generic or medical-specific models including Qwen2-VL-72B, and MedVInT-7B, warranting its efficiency and reliability for real-world deployment.
  3. 3.
    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating explicit thinking processes improves generalization. In medical VQA, we find that RL without reasoning steps yields more accurate predictions, while free-form reasoning—learned from general-domain data—can lead to hallucinations under domain shift. This highlights that in specialized domains—where image types, task goals, and language style differ significantly from general VLM training—the quality of reasoning, rather than its quantity, is critical for generalization.
2 Related Works
General VLMs and Medical VLMs. General-purpose VLMs such as CLIP [15] and BLIP-2 [16] have advanced natural image-text understanding via large-scale pretraining, but struggle with domain-specific tasks like medical reasoning. Recent efforts adapt VLMs through supervised fine-tuning (SFT) on medical datasets, as seen in LLaVA-Med [8] and Med-Flamingo [6]. While effective in-domain, these models often overfit to narrow corpora and lack generalization across modalities or task types. Our work addresses this limitation by introducing RL for scalable, modality-agnostic adaptation.
Reinforcement Learning for Post-Training. RL has shown promise for aligning language models with desired behavior via reward feedback [17, 18]. In vision-language tasks, RL improves VQA accuracy [11] and reduces hallucination [12], but often relies on complex reward models or costly human supervision. GRPO [12] offers a scalable alternative by using rule-based rewards and group-relative comparisons. We extend this approach to medical VQA, enabling efficient adaptation without modality-specific supervision.
Medical Reasoning and Interpretability Interpretable reasoning is critical in medical AI, with recent work exploring CoT prompting [19] and program-guided logic [4]. However, such methods rely on costly expert annotations [20], limiting scalability in clinical domains. Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision. Concurrently, MedVLM-R1 [21] also applies GRPO-based RL to radiology VQA, which is conducted on a relatively small dataset without generalization analysis, making its reasoning robustness challenging to assess. Nonetheless, it highlights the growing interest in RL for medical VLMs.
Figure 2: Medical VQA examples of Med-R1 The left panel shows a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray). The right panel illustrates a disease diagnosis task, where Med-R1 accurately detects urolithiasis in an abdominal CT scan. In both cases, the model provides reasoning (“Think”), explaining its decision-making process before selecting the final answer.
3 Method
We adopt rule-based RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [22] and visual navigation [11]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [23] to train a large base MLLM [24] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the algorithm, our reward design, and data structure below.
3.1 Group Relative Policy Optimization (GRPO)
Overview: RL-based algorithms such as PPO [13] and GRPO [12] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities. Unlike supervised fine-tuning, which directly optimizes maximum likelihood, these RL-based methods instead optimize the policy gradient using reward signals, encouraging reasoning by exploring a much larger solution space. GRPO is closely related to PPO but differs in two key aspects: first, GRPO estimates the advantage using group-based estimation rather than a value function; second, it employs a set of fixed rules as the reward signal instead of a learned reward model. These optimizations make GRPO 50% more resource- and computation-efficient than PPO [12].
Definition: Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration. Let $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta_{\text{new}}}$ denote the old policy and current (new) policy, respectively, where $o$ is a complete response sampled from a policy. Let $\pi_{\theta_{\text{ref}}}$ denote the reference policy, which in practice is the frozen base MLLM. Let $G$ be the number of responses sampled per question in each iteration. The GRPO objective is given by:
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\ \mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)}$
$\quad\frac{1}{G}\sum_{i=1}^{G}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},$
$\qquad\text{clip}\left(\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)A_{i}\Bigg{)}$
$\quad\quad-\beta\,\mathbb{D}_{\text{KL}}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{]}$ (1)
where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage, and $\epsilon$ is the clipping threshold for policy updates The KL divergence term [25] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$. Unlike PPO, which uses a critic model to estimate the advantage $A_{i}$ for a single response $o$, GRPO estimates the relative advantage by sampling a group of responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their rewards within the group to compute a relative advantage [12, 26]. Each reward is calculated based on rules without reward models. We detail reward design below.
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.
RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 98.30 74.43 80.61 51.93 64.17 65.39 70.52 66.13 71.44
MRI 62.33 99.29 73.75 53.18 75.04 64.57 75.59 66.31 71.26
X-Ray 79.67 68.29 95.85 49.32 70.75 76.87 70.28 67.75 72.35
US 55.88 63.96 70.53 99.86 64.78 63.02 63.92 71.53 69.19
Der 53.81 67.36 73.31 49.80 95.10 70.22 64.03 67.93 67.71
FP 54.98 65.95 75.79 49.08 71.36 92.44 65.33 66.40 67.67
OCT 64.48 70.80 77.09 52.65 72.89 73.41 99.17 66.85 72.17
Micro 60.97 63.25 74.92 52.56 64.55 63.66 66.86 93.51 67.54
Overall 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
  • CT - Computed Tomography; MRI - Magnetic Resonance Imaging; US - Ultrasound; Der - Dermoscopy; FP - Fundus Photography.
  • OCT - Optical Coherence Tomography; Micro - Microscopy Images; X-Ray - X-Ray Imaging
Table 2: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.
No-Think RL fine-tuned VLM
Train Test CT MRI X-Ray US Der FP OCT Micro Overall
CT 99.54 82.18 80.00 46.24 72.13 72.50 76.06 70.36 74.88
MRI 79.27 99.72 74.74 42.86 79.94 73.50 84.67 69.10 75.48
X-Ray 78.37 69.94 97.09 46.43 70.75 80.97 69.81 70.63 73.00
US 59.82 66.53 73.37 100.00 70.60 68.03 68.63 79.28 73.28
Der 56.34 67.94 73.87 46.14 98.00 69.67 67.45 67.57 68.37
FP 59.55 68.62 73.93 51.59 72.43 93.99 72.64 70.36 70.39
OCT 80.47 73.50 77.15 49.08 65.08 70.95 100.00 67.21 72.93
Micro 64.15 66.56 76.84 53.13 68.91 70.95 73.47 98.56 71.50
Overall 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.48
Reward design: We follow [12] and use two types of reward: format and accuracy. Firstly, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag. The format reward is designed to check if the aforementioned tags are present in the final response. A reward score of 1 will be given if they exist and are correct. This helps the model to organize its thoughts and answer in a structured format for the ease of reading. The accuracy reward is a rule-based reward that checks if the actual answer matches with the ground truth. Similarly, a reward score of 1 is given when the results match. In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
3.2 No-Thinking-Med-R1
Previous work [14] found that, in image classification tasks, removing both reasoning and format supervision during RL could sometimes improve performance. However, that finding is specific to structured classification tasks with simple output space, where reasoning is rarely required. In contrast, medical VQA involves multimodal, semantically complex inputs, and reasoning failures can arise from domain mismatch rather than verbosity alone [12, 6]. Given the difficulty of obtaining reliable CoT annotations in medical settings, we further investigate the role of intermediate reasoning in RL post-training. We revise the instruction prompt as {Question}. Output the single-letter choice (A, B, C, D,...) in <answer>...</answer> tags., where {Question} will be replaced by each specific question. By doing so, we encourage the model only to output the final answer without any explicit thinking process. The accuracy reward is maintained, where the reward score is 1 when the extracted answer matches the ground truth labels. Note that when there is any text outside the <answer> tag, i.e., an explicit thinking process exists, the extracted content will be null, and therefore the accuracy reward will be 0. Therefore, the model will be forced to generate only the answers.
Table 3: Performance comparison across eight medical modalities. Our GRPO-finetuned model outperforms zero-shot general-purpose VLMs, medical-domain VLMs, and supervised fine-tuning baselines, while maintaining scalability. The best and second-best performances per column are highlighted in red and blue. Modality abbreviations: CT – Computed Tomography; MRI – Magnetic Resonance Imaging; US – Ultrasound; Der – Dermoscopy; FP – Fundus Photography; OCT – Optical Coherence Tomography; Micro – Microscopy; X-Ray – X-ray Imaging.
Methods Modality CT MRI X-Ray US Der FP OCT Micro Overall
Zero-shot VLMs
[…]
$\text{InstructBLIP}^{\dagger}$ [27] 28.72 33.15 61.04 41.25 62.22 50.31 42.59 46.29 45.70
$\text{LLaVA}^{\dagger}$ [28] 17.73 26.72 30.70 18.66 49.74 47.11 33.73 28.87 31.66
[…]
Qwen2-VL-2B [3] 45.10 38.57 39.32 30.86 35.83 43.17 35.14 36.85 38.11
Qwen2-VL-7B [3] 61.46 45.77 64.27 36.01 49.08 59.84 59.32 61.08 54.60
Qwen2-VL-72B [30] 67.97 69.39 77.21 51.39 65.31 72.58 72.76 67.83 68.05
Qwen2.5-VL-3B [31] 53.87 54.23 61.84 32.69 52.94 62.47 56.23 59.64 54.24
Qwen2.5-VL-7B [31] 60.44 58.44 73.99 30.66 62.48 67.30 61.20 67.84 60.29
Qwen2.5-VL-72B [31] 66.18 68.74 77.59 49.81 69.75 71.04 69.22 69.37 67.71
Zero-shot Medical VLMs
[…]
$\text{RadFM}^{\dagger}$ [32] 27.56 24.06 30.95 16.57 39.21 36.89 32.80 27.97 29.50
[…]
$\text{MedVInT}^{\dagger}$ [7] 40.74 43.10 55.10 41.26 29.11 31.84 23.26 32.00 37.05
Fine-tuned VLMs
Qwen2-VL-2B (SFT) 51.74 52.83 65.57 47.65 51.91 52.26 53.99 56.58 54.07
Qwen2.5-VL-3B (SFT) 56.06 60.81 69.23 41.77 60.11 69.19 63.95 65.66 60.85
Qwen2-VL-2B (GRPO + Think) 66.30 71.67 77.73 57.31 72.33 71.20 71.96 70.80 69.91
Qwen2-VL-2B (GRPO + Nothink) 72.19 74.37 78.37 54.43 74.73 75.07 76.59 74.13 72.49
Qwen2.5-VL-3B (GRPO + Nothink) 69.89 72.91 84.52 43.91 73.62 80.10 84.18 71.44 72.57
4 Experiment & Results
4.1 Setup
Datasets. We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [33], which consists of a total of $82,059$ images and $88,996$ vision question answering pairs. OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($,64$ 646), and Microscopy ($5,680$). It is also categorized into five VQA question types, including Anatomy Identification ($16,448$), Disease Diagnosis ($55,387$), Lesion Grading ($2,098$), Modality Recognition ($11,565$), and Other Biological Attributes ($3,498$). We split the dataset into training and test sets following an 80-20 ratio for each setting.
Implementation Details Training is conducted on HGX H100 [34] server with 2×H100 GPUs (80GB VRAM) using PyTorch [35] and FlashAttention-2 [36] for optimized efficiency. We initialize from Qwen2-VL-2B-Instruct [3] with full parameter tuning, employing per-GPU batch size 1 (effective batch size 4 via 2-step gradient accumulation) and bfloat16 mixed precision. Input sequences combine visual embeddings from 328×328 resolution images (max 401k pixels) with textual prompts truncated to 1,024 tokens. The GRPO policy generates four candidate rationales per sample, with a sampling temperature of $\tau=0.7$. Each training is run for one epoch.
Task setting. We evaluate our approach in two distinct generalization settings using the OmniMedVQA dataset [33]: cross-modality generalization and cross-task generalization.
  • Cross-modality generalization: We train our model on a single modality at a time (out of 8 available modalities) and evaluate its performance on the other seven modalities.
  • Cross-task generalization: We identify 5 distinct task types within the dataset and adopt the same train-test partitioning strategy as in the cross-modality setting, training on one task and evaluating on the other four.
Baseline Methods & Evaluation Metric. Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [16], InstructBLIP [27], and the Qwen2-VL series (2B/7B/72B) [3, 30]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [6], MedVInT [7], LLaVA-Med [8]); and SFT variants of Qwen2-VL-2B.
We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where models select answers from clinically validated options. Given a medical image $I$, a question $Q$, and $K$ candidate answers $\{A_{k}\}_{k=1}^{K}$ (one correct and $K-1$ distractors), accuracy is computed as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{y}_{i}=y_{i}\right)$ (2)
where $N$ is the total number of test cases, $\hat{y}_{i}$ is the model’s predicted answer index, $y_{i}$ is the ground truth index, and $\mathbb{I}$ is the indicator function, which returns 1 if the predicted answer matches the ground truth and 0 otherwise.
4.2 Cross-Modality Generalization
We comprehensively evaluate Med-R1’s adaptability across eight medical imaging modalities, including Computed Tomography, Magnetic Resonance Imaging, Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography, Microscopy Images, and X-ray Imaging. Our experiments focus on two key aspects: (1) cross-modal generalization, where the model is trained on one modality and tested on another, and (2) comparative performance against other popular VLMs and medical-specific VLMs evaluated using zero-shot and SFT.
Table 4: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition. The base model is Qwen2-VL-2B.
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
Overall 62.88 66.08 65.87 98.24 80.14 74.64
Table 5: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. The base model is Qwen2-VL-2B.
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification 96.33 59.29 64.22 99.04 76.56 79.09
Disease Diagnosis 55.00 98.30 63.76 98.08 85.23 80.07
Lesion Grading 54.22 58.04 96.33 97.37 72.30 75.65
Modality Recognition 56.84 57.95 55.05 99.79 73.15 68.56
Other Attributes 56.29 58.00 52.29 97.90 99.29 72.75
Overall 63.74 66.32 66.33 98.44 81.31 75.22
Results on generalization. To evaluate Med-R1’s cross-modality generalization, we measure its accuracy across eight distinct medical imaging modalities (Table 1). Overall row and column summarize the model’s average performance across training and test domains, providing insights into its generalization ability. Med-R1 achieves a strong overall accuracy of 69.91%, demonstrating its ability to generalize across diverse medical imaging modalities. Notably, models trained on CT, MRI, and X-Ray exhibit the highest generalization capability, with overall scores of 71.44%, 71.26%, and 72.35%, respectively. In contrast, models trained on Fundus Photography and Microscopy images show lower generalization, with 67.67% and 67.54% overall accuracy, indicating that certain modality-specific features (e.g., texture-based imaging in US and Micro) may not transfer as effectively to other domains. Importantly, the overall test accuracy of 69.91% highlights Med-R1’s ability to perform well across unseen imaging modalities, despite being trained on a single domain at a time. This result underscores the effectiveness of reinforcement learning in enhancing cross-modality transfer, allowing the model to maintain robust performance without requiring extensive retraining for each medical imaging modality.
Table 6: Performance Comparison of VLMs on Five Medical VQA Tasks: GRPO Fine-Tuning Outperforms Zero-Shot and SFT Baselines Across Diverse Reasoning Tasks. Performance is evaluated on five clinical reasoning types (columns) across three model categories: general-purpose VLMs (zero-shot), medical VLMs (zero-shot), and fine-tuned VLMs. The best and second-best performances are marked in red and blue.
Methods Types Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Zero-shot VLMs
[…]
$\text{InstructBLIP}^{\dagger}$ [27] 44.35 32.29 59.25 75.27 23.72 40.40
$\text{LLaVA}^{\dagger}$ [28] 25.86 29.10 43.95 21.36 31.90 27.96
[…]
Qwen2-VL-2B [3] 30.70 36.53 43.58 59.90 42.19 42.58
Qwen2-VL-7B [3] 42.57 48.83 52.06 84.74 59.66 57.57
Qwen2-VL-72B [30] 56.41 65.71 62.15 98.11 80.53 72.58
Qwen2.5-VL-3B [31] 35.23 50.45 52.79 85.23 54.77 55.69
Qwen2.5-VL-7B [31] 41.00 61.32 54.13 97.78 70.45 64.94
Qwen2.5-VL-72B [31] 57.22 62.55 60.32 98.20 77.41 71.14
Zero-shot Medical VLMs
[…]
$\text{RadFM}^{\dagger}$ [32] 13.31 21.69 30.35 26.64 43.85 26.99
[…]
$\text{MedVInT}^{\dagger}$ [7] 40.26 35.78 12.77 68.10 30.30 40.04
Fine-tuned VLMs
Qwen2-VL-2B (SFT) 53.97 51.62 60.71 86.77 63.91 63.39
Qwen2.5-VL-3B (SFT) 54.91 57.75 64.04 84.95 71.56 66.64
Qwen2-VL-2B (GRPO + Think) 62.88 66.08 65.87 98.24 80.14 74.64
Qwen2-VL-2B (GRPO + Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
Qwen2.5-VL-3B (GRPO + Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
Comparison to zero-shot and SFT evaluations with other VLMs. As demonstrated in Table 3, Med-R1 demonstrates its superiority across all eight medical imaging modalities while maintaining exceptional parameter efficiency. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular modality. For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities. Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the 36×36$\times$36 × parameter disparity. This advantage amplifies in critical diagnostic tasks: Med-R1 attains 71.67% accuracy in MRI compared to Qwen2-VL-72B’s 69.39%, and achieves 72.33% versus 65.31% in dermoscopy, challenging the prevailing scale-equals-performance paradigm. The limitations of specialized medical VLMs become evident through Med-Flamingo’s 30.38% average accuracy, which Med-R1 outperforms by 39.53%. This stark contrast underscores the ineffectiveness of narrow medical pretraining compared to our RL-driven adaptation strategy. When compared to supervised fine-tuning approaches, GRPO delivers 15.84% accuracy gains over SFT-tuned Qwen2-VL-2B (69.91% vs. 54.07%), with particularly significant improvements in CT interpretation (66.30% vs. 51.74%) and OCT analysis (71.96% vs. 53.99%).
4.3 Cross-Task Generalization
We also evaluate Med-R1’s generalization across five important clinical tasks [33]: Anatomy Identification, Disease Diagnosis, Lesion Grading, Modality Recognition, and Other Attributes. Similar to subsection 4.2, we focus our evaluation on two aspects: cross-modality generalization and comparison against SFT and zero-shot with other VLMs.
Figure 3: Cross-modality accuracy difference between No-Thinking-Med-R1 and Med-R1. Each cell shows the performance gap (%) when trained on one modality (y-axis) and evaluated on others (x-axis). Red indicates improvement; blue indicates degradation.
Figure 4: Cross-task accuracy difference between No-Thinking-Med-R1 and Med-R1. Each cell shows the performance gap (%) when trained on one task (y-axis) and tested on another (x-axis). Red indicates improvement; blue indicates degradation.
Results on generalization. As shown in Table 4, models trained on “disease diagnosis" data achieve the best generalization, with 81.64% overall accuracy. This suggests that disease diagnosis encompasses diverse feature representations that transfer well across tasks, likely due to its reliance on both anatomical and pathological cues. In contrast, models trained on “modality recognition" exhibit strong generalization in task-agnostic settings (98.24% in the “modality recognition" column), indicating that learning modality distinctions aids in extracting transferable image features. However, training on “lesion grading" leads to high in-task performance (86.24%) but relatively lower transferability, implying that this task captures more specialized features that do not generalize as effectively. These results highlight the trade-off between specialization and adaptability, emphasizing the importance of task selection when designing models for broad medical applications.
Comparison to zero-shot and SFT evaluations with other VLMs. Table 6 shows the comparison results with other popular VLMs evaluated with zero-shot and SFT. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular task. For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given task when evaluated using models that were separately fine-tuned on each of the five training tasks. First of all, the results clearly show that Med-R1 outperforms all other popular VLMs’ zero-shot generalization. Remarkably, Med-R1 even outperforms Qwen2-VL-72B (74.64% vs. 72.58%), a model with 70 billion more parameters. More importantly, this suggests that RL can effectively elevate small models with moderate capacity, opening doors for many real-world applications where resource is a constraint. In contrast, the average generalization with the identical base model trained with SFT is merely 63.39%, 11.25% below Med-R1, further demonstrating the strong generalization capability of Med-R1.
Figure 5: Rewards & Length. The top row shows the change in RL rewards during Med-R1 training across eight imaging modalities and five task types. We observe that the model quickly converges to stable reward values within 100–200 steps, indicating the computational efficiency of our RL post-training process. The bottom row shows the output completion length over training steps. Interestingly, the output length gradually decreases as training progresses, suggesting that the reward function discourages overly verbose or unfaithful reasoning. This observation challenges the conventional assumption that longer reasoning necessarily benefits vision-language performance in VQA.
4.4 Analysis of Thinking Process
We present the results of No-Thinking-Med-R1 in this section. Cross-task performance is shown in Table 5, and cross-modality performance is shown in Table 2. Overall, No-Thinking-Med-R1 achieves stronger in-domain performance across all settings compared to Med-R1, suggesting that removing explicit reasoning generation may lead to more effective task-specific learning. To assess generalization, we compare Med-R1 and No-Thinking-Med-R1 across modalities and tasks. As visualized in Figure 3 and Figure 4, No-Thinking-Med-R1 consistently outperforms Med-R1 in cross-modality settings. In cross-task scenarios, however, the results are more mixed—showing gains in some tasks and declines in others. These findings reinforce the practical motivation behind Med-R1. In clinical domains, acquiring high-quality CoT supervision is prohibitively costly and often impractical. While reasoning improves performance in general domains, our results show that this does not necessarily hold in medical settings, where unsupervised rationales may become unreliable under domain shift. Med-R1 demonstrates that even without reasoning supervision, RL can offer a robust and efficient path for adapting VLMs to medicine.
4.5 Analysis of Rewards and Completion Length.
The top row of Figure 5 shows the reward dynamics during GRPO-based reinforcement learning. Across all modalities and tasks, Med-R1 converges within 100–200 steps, as evidenced by the log-scale reward curves. This rapid and consistent convergence highlights the efficiency of our training pipeline and the model’s ability to quickly adapt to the reward structure. Moreover, the smooth reward trajectories suggest that GRPO remains stable even with limited RL updates, making it well-suited for low-resource scenarios.
Unlike in scientific domains such as mathematics [12], where model responses become increasingly detailed as thinking deepens, we observe a different trend in Med-RL training: the average response length decreases rapidly and then fluctuates (Figure 5). Rather than generating elaborate rationales, the model learns to produce shorter, more direct outputs—potentially favoring conciseness over verbosity. This aligns with our findings in subsection 4.4, suggesting that extended reasoning may not always benefit cross-domain VQA, especially under large domain shifts. We hypothesize that RL without CoT supervision encourages the model to rely on its pre-trained knowledge when generating rationales. While this may be effective in general domains, it can lead to hallucinated or misaligned reasoning in specialized domains like medical images. These results challenge the common assumption that more thinking always improves performance, and highlight that the “quality”—rather than the “presence”—of thinking is a crucial factor affecting the outcome of RL post-training.
4.6 Limitations and Future Work.
This work marks an initial step in applying RL to medical vision-language models. We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging. In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time. Future directions include extending Med-R1 to support multiple inputs. We also plan to incorporate patient context and explore more advanced reasoning frameworks for clinical deployment.
5 Conclusion
We present Med-R1, a RL-enhanced vision-language model that improves medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 surpasses the limitations of supervised fine-tuning, achieving strong cross-modality and cross-task generalization. Despite its compact 2B size, it matches or exceeds much larger models—including domain-specific and general-purpose VLMs—while remaining lightweight and deployable in resource-constrained clinical settings. Interestingly, we find that RL-based learning without explicit step-by-step rationales can outperform reasoning-heavy approaches, particularly under domain shift. This suggests that the quality, not quantity, of reasoning is key for robust medical adaptation. Med-R1 opens promising directions for developing scalable, reliable, and efficient AI systems in healthcare.